Hi i am studying for an algebra placement exam on Monday. I used the linear system solver for
3x+4y = 27
-3x+4y = -27
I get the substitution as far as x a function of y from the first equation
x = (27 - 4y)/3
But why does the -4y become +4y when you substitute it into the second equation?
-3(27+4y)/3+4y=-27
Also what are the steps in simplifying that takes you from -3(27+4y)/3+4y=-27
to: 24y=0
or y=0
Trying to comprehend!

Answer by wuwei96815(245)

Are you required to use substitution to solve this set of equations? There is a simpler method. Just multiply the first equation by -1 and add the equations together.

-1(3x+4y=27) becomes -3x-4y=-27


-3x-4y=-27 plus -3x+4y=-27 becomes -6x=-54


x=9

Now, substitute the value for x in either equation to find y:
3x+4y=27
3(9)+4y=27
y=0

You can verify that x=9 and y=0 by substituting these values into any of the equations.